# Finite-size corrections to the energy spectra of gapless one-dimensional systems in the presence of boundaries

### v1

@@ 1 Introduction
We note that, although BCFT has been studied in various topics [25, 26, 27], a detailed systematic study of the finite-size spectrum in the general context is lacking. For the TF-Ising model, there exist some articles that relate finite-size corrections to conformal perturbation theory instead of the boundary one [28, 29, 30, 31]. Although they provide correct universal ratios, we will show that such results are only correct for specific operators. Working in the context of BCFT is indispensable for generic perturbation in other models.
@@ 2.1.1 General aspect
We find the perturbative corrections as follows:
@@
$\left(\delta\mathcal{E}^{\alpha\beta}_{0}\right)^{(2)}_{\text{boundary}}$ $=\sum_{j}\left(\delta\mathcal{E}^{\alpha\beta}_{0}\right)^{(2)}_{g^{L}_{j},g^{R}_{j}}\ ,$ (11)
@@
$\left(\delta\mathcal{E}^{\alpha\beta}_{0}\right)^{(2)}_{g^{L}_{j},g^{R}_{j}}$ $\sim-\left(\frac{\pi}{L}\right)^{2h_{j}-1}\left[\left({g^{L}_{j}}^{2}+{g^{R}_{j}}^{2}\right)\frac{\Gamma(1/2-h_{j})\Gamma(h_{j})}{2^{2h_{j}}\sqrt{\pi}}+\frac{2g^{R}_{j}g^{L}_{j}}{h_{j}}{}_{2}F_{1}(h_{j};2h_{j};1+h_{j};-1)\right]$ (12)
[…]
@@
$\left(\delta\mathcal{E}^{\alpha\beta}_{0}\right)^{(2)}_{g^{L}_{j},g^{R}_{j}}$ $\sim-\left[\left({g^{L}_{j}}^{2}+{g^{R}_{j}}^{2}\right)\left(\text{const.}+\log{\frac{L}{a}}\right)+2g^{R}_{j}g^{L}_{j}\frac{\pi}{2}\right]\ ,$ (13)
@@ 2.1.2 Energy-momentum tensor perturbations
Any CFT has the boundary operator of scaling dimension $2$, which can be identified with the (holomorphic part of) the energy-momentum tensor $T$. For the boundary perturbation $T$, the first-order correction to the energy eigenvalues is proportional to
$g^{L,R}_{T}\left(\frac{\pi}{L}\right)^{2}\ ,$ (14)
since $h_{T}=2$. However, this correction can be absorbed by renormalizing the system size at least in the first order in $g^{L,R}_{T}$, since
$\frac{1}{L+\delta L}\sim\frac{1}{L}-\frac{\delta L}{L^{2}}+O(\frac{1}{L^{3}})\ .$ (15)
Thus, as long as we use the appropriate effective length for a lattice model, we do not have to consider this perturbation explicitly.
@@
$L=Na+\left(\delta L\right)_{L}+\left(\delta L\right)_{R}\ ,$ (16)
where $\left(\delta L\right)_{L,R}$ are the non-universal corrections of order of $a$, determined by the boundary conditions on each end of the lattice model.
@@
$g_{T}\int\mathrm{d}x\;\left[T(x)+\bar{T}(x)\right]\ .$ (17)
@@
$g_{T^{2}}\int\mathrm{d}x\;\left[T^{2}(x)+\bar{T}^{2}(x)\right]\ ,$ (18)
which has the conformal dimension $(4,0)$ and $(0,4)$, is generally present [36]. Although it is a renormalization group-irrelevant perturbation, it does have nontrivial effects on energy levels:
@@
$A(0,r)$ $=\left(\frac{11}{30}+\frac{c}{12}\right)r\left(2r^{2}-3\right)\quad(r\neq 1)\ ,$ (21)
$A(h,r)$ $=(h+r)\left(\left(h-\frac{1}{6}-\frac{c}{12}\right)+\frac{r(2h+r)(5h+1)}{(h+1)(2h+1)}\right)\quad h\neq 0\ .$ (22)
@@ 2.2.1 Free boundary ($\zeta_{R}=\zeta_{L}=\zeta=0$)
Namely, the corrections to the length for the free boundary ($\zeta=0$) is
$\left(\delta L\right)_{\text{free}}=+\frac{1}{4}a\ .$ (27)
@@
$E_{2}(\text{free},\text{free})$ $=E_{0}(\text{free},\text{free})+\frac{\pi 3v}{2L}-\frac{9v}{64}\left(\frac{\pi}{L}\right)^{3}+\cdots\ ,$ (29)
$E_{3}(\text{free},\text{free})$ $=E_{0}(\text{free},\text{free})+\frac{\pi 4v}{2L}-\frac{7v}{48}\left(\frac{\pi}{L}\right)^{3}+\cdots\ ,$ (30)
@@
$\left(\delta\mathcal{E}_{2}\right)^{(1)}_{T^{2}}-\left(\delta\mathcal{E}_{0}\right)^{(1)}_{T^{2}}$ =$1/2,1|H_{T^{2}}|1/2,1\right>-\left<0,0|H_{T^{2}}|0,0\right>=2gv\frac{\pi^{4}}{L^{3}}\frac{63}{16}\ ,$ (33)
$\left(\delta\mathcal{E}_{3}\right)^{(1)}_{T^{2}}-\left(\delta\mathcal{E}_{0}\right)^{(1)}_{T^{2}}$ =$0,2|H_{T^{2}}|0,2\right>-\left<0,0|H_{T^{2}}|0,0\right>=2gv\frac{\pi^{4}}{L^{3}}\frac{49}{12}\ ,$ (34)
@@ 2.2.2 Fixed boundary conditions
$\left(\delta L\right)_{\uparrow/\downarrow}=-\frac{1}{4}a\ .$ (39)
@@
$E_{0}(\uparrow/\downarrow,\text{free})$ $=2\left(\frac{1}{2}\right)-\frac{4}{\pi}N-\frac{\pi}{12N}-\frac{1}{2880}\left(\frac{\pi}{N}\right)^{3}+\cdots\ .$ (40)
@@
$L=Na+(\delta L)_{\uparrow/\downarrow}+(\delta L)_{\text{free}}=Na\ .$ (41)
@@
$\left(\delta\mathcal{E}_{0}\right)^{(1)}_{T^{2}}$ $=\langle 1/16,0|H_{T^{2}}|1/16,0\rangle=2g_{T^{2}}v\frac{\pi^{4}}{L^{3}}\left(-\frac{7}{1440}\right)\ ,$ (42)
@@ 2.2.3 Boundary field perturbation to the free boundary
$\left(\delta\mathcal{E}_{0}\right)^{(2)}_{\zeta}$ $=-2v{c_{\sigma}}^{2}\zeta^{2}\left\{\text{const.}+\log{\frac{L}{a}}\pm\frac{\pi}{2}\right\}\ ,$ (45)
@@ 2.3 Application to the three-state Potts model
where:
$M=\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ 1&0&0\end{array}\right);\quad R=\left(\begin{array}[]{ccc}e^{2\pi i/3}&0&0\\ 0&e^{4\pi i/3}&0\\ 0&0&1\end{array}\right)\ .$ (56)
@@ 2.3.2 Fixed/Mixed boundary conditions
$E_{0}(N+1,\text{A},\text{A})=E_{0}(N+1,\text{B},\text{B})=E_{0}(N+1,\text{C},\text{C})$ $=E_{0}(N,\zeta_{L}=0,\zeta_{R}=0)\ ,$ (63)
@@
$\left(\delta L\right)_{\text{A,B,C}}=\left(\delta L\right)_{\text{free}}-\frac{1}{2}a.$ (67)
@@
$\left(\delta L\right)_{\text{A,B,C}}=-\frac{1}{2}a,$ (68)
@@
$M_{\mathrm{AB}}=\begin{pmatrix}0&0&0\\ 0&0&1\\ 0&0&0\\ \end{pmatrix},\quad M_{\mathrm{BC}}=\begin{pmatrix}0&1&0\\ 0&0&0\\ 0&0&0\\ \end{pmatrix}.,$
@@
$(\delta L)_{\text{AB,BC,CA}}=(\delta L)_{\text{A,B,C}}=-\frac{1}{2}a\ .$ (74)
@@ 3.2 Boundary conformal field theory: a review
where $w=x+iy$ [^2] For simplicity, when we are not referring to any particular lattice model, we set $v=1$ in the field theory result.. We can map this infinite long strip onto the upper half of the complex plane through a conformal mapping
@@ 3.4 Finite-size corrections from bulk $T$
We also note that the Lorentz invariance-breaking marginal bulk perturbation $T+\bar{T}$ is also allowed in lattice models. However, it just corresponds to the renormalization of the spin-wave velocity (“speed of light”).
@@
Furthermore, we can also consider the $T\bar{T}$ perturbation [36],
$H_{T\bar{T}}$ $=g\int_{0}^{L}T(w)\bar{T}(\bar{w})\,\mathrm{d}y$
$=2\pi g\left[\frac{1}{2\pi i}\int_{C^{+}}\left(\frac{\pi}{L}\right)^{3}\left(z^{2}\mathbb{T}(z)-\frac{c}{24}\right)\left(z^{*2}\mathbb{T}(z^{*})-\frac{c}{24}\right)\frac{\mathrm{d}z}{z}\right]$
$=gL\left(\frac{\pi}{L}\right)^{4}\left[\sum_{n\in\mathbb{Z}}L^{\text{(H)}}_{n}L_{n}^{\text{(H)}}-\frac{c}{12}L^{\text{(H)}}_{0}+\left(\frac{c}{24}\right)^{2}\right]$ (96)
@@
$\delta\left(\mathcal{E}_{n}\right)_{T\bar{T}}^{\text{(1)}}=\left\langle\Delta,r\middle|H_{T\bar{T}}\middle|\Delta,r\right\rangle=\pi g\frac{\pi^{3}}{L^{3}}\left(\Delta+r-\frac{c}{24}\right)^{2}\ .$ (97)
However, for more general bulk perturbations, the treatment of the boundary is more complicated and generally depends on each conformally invariant boundary condition.
@@ 3.5 UV divergence and regularization
The first term is singular when $h_{\Psi}\geq\frac{1}{2}$. To regularise it, we introduce a UV-cutoff, namely the lattice spacing $a$,
$\int_{a}^{\infty}\frac{\mathrm{d}\tau}{\left[2\sinh(\frac{\pi\tau}{2L})\right]^{2h_{\Psi}}}=\left(\frac{L}{\pi}\right)\left\{\frac{\Gamma(1/2-h_{\Psi})\Gamma(h_{\Psi})}{2^{2h_{\Psi}}\sqrt{\pi}}+\left(\frac{L}{\pi a}\right)^{2h_{\Psi}-1}\left[\frac{\Gamma(1/2-h_{\Psi})}{2\Gamma(3/2-h_{\Psi})}+O(\frac{a}{L})\right]\right\}\ .$ (100)
@@
$\int_{a}^{\infty}\frac{\mathrm{d}\tau}{2\sinh(\frac{\pi\tau}{2L})}=\left(\frac{L}{\pi}\right)\left[2\log 2-\log\left(\frac{\pi a}{L}\right)+O(\frac{a}{L})\right].$ (101)

### v2

@@ 1 Introduction
We note that, although BCFT has been studied in various topics [25, 26, 27, 28], a detailed systematic study of the finite-size spectrum in the general context is lacking. For the TF-Ising model, there exist some articles that relate finite-size corrections to conformal perturbation theory instead of the boundary one [29, 30, 31, 32]. Although they provide correct universal ratios, we will show that such results are only correct for specific operators. Working in the context of BCFT is indispensable for generic perturbation in other models.
@@ 2.1.1 General aspect
We find the perturbative corrections as follows [^1] For most of the discussions in this paper, we assume that the bulk/boundary operators are primaries just for simplicity. One can easily obtain the results for descendants by applying Virasoro operators to the correlation functions.:
@@
$\left(\delta\mathcal{E}^{\alpha\alpha}_{0}\right)^{(2)}_{\text{boundary}}$ $=\sum_{j}\left(\delta\mathcal{E}^{\alpha\alpha}_{0}\right)^{(2)}_{g^{L}_{j},g^{R}_{j}}\ ,$ (11)
@@
$\left(\delta\mathcal{E}^{\alpha\alpha}_{0}\right)^{(2)}_{g^{L}_{j},g^{R}_{j}}$ $\sim-\left(\frac{\pi}{L}\right)^{2h_{j}-1}\left[\left({g^{L}_{j}}^{2}+{g^{R}_{j}}^{2}\right)\frac{\Gamma(1/2-h_{j})\Gamma(h_{j})}{2^{2h_{j}}\sqrt{\pi}}+\frac{2g^{R}_{j}g^{L}_{j}}{h_{j}}{}_{2}F_{1}(h_{j};2h_{j};1+h_{j};-1)\right]$ (12)
$\quad+\frac{\text{const.}}{a^{2h_{j}-1}}\left({g^{L}_{j}}^{2}+{g^{R}_{j}}^{2}\right)\ ,$
@@
$\left(\delta\mathcal{E}^{\alpha\alpha}_{0}\right)^{(2)}_{g^{L}_{j},g^{R}_{j}}$ $\sim-\left[\left({g^{L}_{j}}^{2}+{g^{R}_{j}}^{2}\right)\left(\text{const.}+\log{\frac{L}{a}}\right)+2g^{R}_{j}g^{L}_{j}\frac{\pi}{2}\right]\ ,$ (13)
@@ 2.1.2 Energy-momentum tensor perturbations
Any CFT has the boundary operator of scaling dimension $2$, which can be identified with the (holomorphic part of) the energy-momentum tensor $T$. For the boundary perturbation $T$, the first-order correction to the energy eigenvalues is
$\left(\delta\mathcal{E}_{n}^{\alpha\beta}\right)_{T}=\left(\frac{\pi}{L}\right)^{2}(g^{\alpha}_{T}+g^{\beta}_{T})\left(h_{n}-\frac{c}{24}\right),$ (14)
where $g_{T}^{\alpha/\beta}$ is the coupling constant of $T$ associated with boundary condition $\alpha/\beta$. However, this correction can be absorbed by renormalizing the system size at least in the first order in $g_{T}$, since
$\mathcal{E}^{\alpha\beta}_{n}(L)+\left(\delta\mathcal{E}_{n}^{\alpha\beta}\right)_{T}(L)$ $=\left(\frac{\pi}{L}\right)\left[1+(g^{\alpha}_{T}+g^{\beta}_{T})\left(\frac{\pi}{L}\right)\right]\left(h_{n}-\frac{c}{24}\right)$
$\sim\frac{\pi}{L+(\delta L)_{\alpha}+\delta(L)_{\beta}}\left(h_{n}-\frac{c}{24}\right)$
$=\mathcal{E}^{\alpha\beta}_{n}(L+(\delta L)_{\alpha}+(\delta L)_{\beta}),$ (15)
where we set the local shift of system size $(\delta L)_{\alpha/\beta}=-\pi g_{T}^{\alpha/\beta}$. Thus, as long as we use the appropriate effective length for a lattice model, we do not have to consider this perturbation explicitly for all energy levels.
@@
$L=\left[N+\left(\delta L\right)_{L}+\left(\delta L\right)_{R}\right]a\ ,$ (16)
where $\left(\delta L\right)_{L,R}$ are the non-universal corrections of order of $O(1)$, determined by the boundary conditions on each end of the lattice model.
@@
$g_{T}\int\mathrm{d}y\;\left[T(w)+\bar{T}(\bar{w})\right]\ .$ (17)
@@
$g_{T^{2}}\int\mathrm{d}y\;\left[T^{2}(w)+\bar{T}^{2}(\bar{w})\right]\ ,$ (18)
which has the conformal dimension $(4,0)$ and $(0,4)$, is generally present in the effective theory of a lattice model[37], since it corresponds to the the breaking of the Lorentz invariance, or equivalently continuous rotation symmetry in the Euclidean space-time.
Although it is a renormalization group-irrelevant perturbation, it does have nontrivial effects on energy levels:
@@
$A(0,r)$ $=\left(\frac{11}{30}+\frac{c}{12}\right)r\left(2r^{2}-3\right),\quad r\neq 1\ ,$
$A(h,r)$ $=(h+r)\left[\left(h-\frac{1}{6}-\frac{c}{12}\right)+\frac{r(2h+r)(5h+1)}{(h+1)(2h+1)}\right],\quad h\neq 0\ .$ (21)
For completeness, we also considered the $T\bar{T}$ bulk perturbation:
$g_{T\bar{T}}\int\mathrm{d}yT\bar{T}(w,\bar{w})$ (22)
Such perturbation has conformal dimension $(2,2)$, thus at the 1st order giving rise to FSC as
$\delta\left(\mathcal{E}_{n}\right)_{T\bar{T}}^{\text{(1)}}=\pi g_{T\bar{T}}\left(\frac{\pi}{L}\right)^{3}\left(\Delta+r-\frac{c}{24}\right)^{2}\ ,$ (23)
which is at the same order as $T^{2}+\bar{T}^{2}$. For the XXZ chain, these two perturbation together give rise to the the leading-order or next-to-leading-order FSC[38, 39, 40]. However, it does not present in the particular models we are considering in this paper (TF-Ising and three-state Potts)[37].
@@ 2.2.1 Free boundary ($\zeta_{R}=\zeta_{L}=\zeta=0$)
In this paper, to compare the CFT result with the lattice model, we employed the convention that the lattice spacing $a$ is taken to $1$ for simplicity in the CFT expression [^2] It should be noted that to recover the correct dimension in the continuum limit, one requires that $L=\left[N+(\delta L)_{L}+(\delta L)_{R}\right]a$, where $a$ has the dimension of length. See Sec. 4.2 for an example.. By Eq. (16), the corrections to the length for the free boundary ($\zeta=0$) is
$\left(\delta L\right)_{\text{free}}=+\frac{1}{4}\ .$ (28)
@@
$E_{2}(\text{free},\text{free})$ $=E_{0}(\text{free},\text{free})+\frac{3\pi v}{2L}-\frac{9v}{64}\left(\frac{\pi}{L}\right)^{3}+\cdots\ ,$ (30)
$E_{3}(\text{free},\text{free})$ $=E_{0}(\text{free},\text{free})+\frac{2\pi v}{L}-\frac{7v}{48}\left(\frac{\pi}{L}\right)^{3}+\cdots\ ,$ (31)
@@
$\left(\delta\mathcal{E}_{2}\right)^{(1)}_{T^{2}}-\left(\delta\mathcal{E}_{0}\right)^{(1)}_{T^{2}}$ =$1/2,1|H_{T^{2}}|1/2,1\right>-\left<0,0|H_{T^{2}}|0,0\right>=2g_{T^{2}}v\frac{\pi^{4}}{L^{3}}\frac{63}{16}\ ,$ (34)
$\left(\delta\mathcal{E}_{3}\right)^{(1)}_{T^{2}}-\left(\delta\mathcal{E}_{0}\right)^{(1)}_{T^{2}}$ =$0,2|H_{T^{2}}|0,2\right>-\left<0,0|H_{T^{2}}|0,0\right>=2g_{T^{2}}v\frac{\pi^{4}}{L^{3}}\frac{49}{12}\ ,$ (35)
@@ 2.2.2 Fixed boundary conditions
$\left(\delta L\right)_{\uparrow/\downarrow}=-\frac{1}{4}\ .$ (40)
@@
$E_{0}(\uparrow/\downarrow,\text{free})$ $=2\left(\frac{1}{2}\right)-\frac{4}{\pi}N+\frac{\pi}{12N}+\frac{1}{2880}\left(\frac{\pi}{N}\right)^{3}+\cdots\ .$ (41)
@@
$L=N+(\delta L)_{\uparrow/\downarrow}+(\delta L)_{\text{free}}=N\ .$ (42)
@@
$\left(\delta\mathcal{E}_{0}\right)^{(1)}_{T^{2}}$ $=\langle\frac{1}{16},0|H_{T^{2}}|\frac{1}{16},0\rangle=2g_{T^{2}}v\frac{\pi^{4}}{L^{3}}\left(-\frac{7}{1440}\right)\ ,$ (43)
@@ 2.2.3 Boundary field perturbation to the free boundary
$\left(\delta\mathcal{E}_{0}\right)^{(2)}_{\zeta}$ $=-2v{c_{\sigma}}^{2}\zeta^{2}\left\{\text{const.}+\log{L}\pm\frac{\pi}{2}\right\}\ ,$ (46)
@@ 2.3 Application to the three-state Potts model
where
$M=\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ 1&0&0\end{array}\right),\quad R=\left(\begin{array}[]{ccc}e^{2\pi i/3}&0&0\\ 0&e^{4\pi i/3}&0\\ 0&0&1\end{array}\right)\ .$ (57)
@@ 2.3.2 Fixed/Mixed boundary conditions
$E_{0}(N+1,\text{A},\text{A})=E_{0}(N+1,\text{B},\text{B})=E_{0}(N+1,\text{C},\text{C})$ $=E_{0}(N,\text{free},\text{free})\ ,$ (64)
@@
$\left(\delta L\right)_{\text{A,B,C}}=\left(\delta L\right)_{\text{free}}-\frac{1}{2}.$ (68)
@@
$\left(\delta L\right)_{\text{A,B,C}}=-\frac{1}{2},$ (69)
@@
$M_{\mathrm{AB}}=\begin{pmatrix}0&0&0\\ 0&0&1\\ 0&0&0\\ \end{pmatrix},\quad M_{\mathrm{BC}}=\begin{pmatrix}0&1&0\\ 0&0&0\\ 0&0&0\\ \end{pmatrix}.$
@@
$(\delta L)_{\text{AB,BC,CA}}=(\delta L)_{\text{A,B,C}}=-\frac{1}{2}\ .$ (75)
@@ 3.2 Boundary conformal field theory: a review
where $w=x+iy$ [^4] For simplicity of the expressions, when we are not referring to any particular lattice model, we set $v=1$ in the field theory result.. We can map this infinite long strip onto the upper half of the complex plane through a conformal mapping
@@ 3.3 Finite-size corrections from perturbation theory
For decedents, boundary $T$, for instance, will have a first-order perturbation as:
$\delta\left(\mathcal{E}_{n}\right)_{T}^{\text{(1)}}$ $=\left(\frac{\pi}{L}\right)^{2}\big{\langle}\Psi_{n}^{\alpha\beta}\big{|}g_{T}^{L}\left(\mathbb{T}(1)-\frac{c}{24}\right)+g_{T}^{R}\left(\mathbb{T}(-1)-\frac{c}{24}\right)\big{|}\Psi_{n}^{\alpha\beta}\big{\rangle}$
$=(g_{T}^{L}+g_{T}^{R})\left(\frac{\pi}{L}\right)^{2}\left(h_{n}-\frac{c}{24}\right)$ (91)
@@ 3.4 Finite-size corrections from bulk $T$
The Lorentz invariance-breaking marginal bulk perturbation $T+\bar{T}$ is also allowed in lattice models. With the help of the above results, it just gives rise to an renormalization of the spin-wave velocity (“speed of light”), like the PBC case.
One can also consider the $T\bar{T}$ perturbation. This time, the image method fails to apply, but one can still perform the mode expansion to obtain
$H_{T\bar{T}}$ $=g_{T\bar{T}}\int_{0}^{L}T(w)\bar{T}(\bar{w})\,\mathrm{d}y$
$=2\pi g_{T\bar{T}}\left[\frac{1}{2\pi i}\int_{C^{+}}\left(\frac{\pi}{L}\right)^{3}\left(z^{2}\mathbb{T}(z)-\frac{c}{24}\right)\left(z^{*2}\mathbb{T}(z^{*})-\frac{c}{24}\right)\frac{\mathrm{d}z}{z}\right]$
$=\pi g_{T\bar{T}}\left(\frac{\pi}{L}\right)^{3}\left[\sum_{n\in\mathbb{Z}}L^{\text{(H)}}_{n}L_{n}^{\text{(H)}}-\frac{c}{12}L^{\text{(H)}}_{0}+\left(\frac{c}{24}\right)^{2}\right]$ (98)
@@
$\delta\left(\mathcal{E}_{n}\right)_{T\bar{T}}^{\text{(1)}}=\left\langle\Delta,r\middle|H_{T\bar{T}}\middle|\Delta,r\right\rangle=\pi g_{T\bar{T}}\left(\frac{\pi}{L}\right)^{3}\left(\Delta+r-\frac{c}{24}\right)^{2}\ .$ (99)
Such FSC, if it presents, is at the same order as the 1st order perturbation from $T^{2}+\bar{T}^{2}$.
@@ 3.5 UV divergence and regularization
The first term is singular when $h_{\Psi}\geq\frac{1}{2}$. To regularise it, we introduce a UV-cutoff $b$, which can be identified as the discreteness nature of the lattice model, so $b\sim a$
$\int_{b}^{\infty}\frac{\mathrm{d}\tau}{\left[2\sinh(\frac{\pi\tau}{2L})\right]^{2h_{\Psi}}}=\left(\frac{L}{\pi}\right)\left\{\frac{\Gamma(1/2-h_{\Psi})\Gamma(h_{\Psi})}{2^{2h_{\Psi}}\sqrt{\pi}}+\left(\frac{L}{\pi b}\right)^{2h_{\Psi}-1}\left[\frac{\Gamma(1/2-h_{\Psi})}{2\Gamma(3/2-h_{\Psi})}+O\left(\frac{b}{L}\right)\right]\right\}\ .$ (102)
@@
$\int_{a}^{\infty}\frac{\mathrm{d}\tau}{2\sinh(\frac{\pi\tau}{2L})}=\left(\frac{L}{\pi}\right)\left[2\log 2-\log\left(\frac{\pi b}{L}\right)+O\left(\frac{b}{L}\right)\right].$ (103)